OpenStudy (anonymous):

A line has a y-intercept of -6 and a slope of . Which of the following points is located on the line?

OpenStudy (anonymous):

thats a real nice slope ya got there ;)

OpenStudy (ddcamp):

y = mx + b; m is the slope, b is the y intercept. Plug in the x value for x and the y value for y. If the point is on the line, both sides will be equal.

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

-6 and a slope of 3/5*

OpenStudy (ddcamp):

y = (3/5)x - 6

OpenStudy (anonymous):

so your formula is \(y={3\over5}x-6\) plug in the points into the equation and see which set is true

OpenStudy (anonymous):

idk how to do it :/

OpenStudy (ddcamp):

You have some points (a,b), right? See if: b = (3/5)a - 6 is true.

OpenStudy (anonymous):

i still dont understand :/ its my weakness

OpenStudy (ddcamp):

Does the question give you some coordinates to check?

OpenStudy (anonymous):

ya its 1-7

OpenStudy (anonymous):

what are the answer choices?

OpenStudy (anonymous):

A. (-3, -1)B. (-5, -3)C. (5, -3)D. (3, -1)

OpenStudy (ddcamp):

I'll show you how to check A, then see if you can do the others.

OpenStudy (ddcamp):